A siren produces a sound intensity level of 90 decibel at a distance of 8.0 m.


assume that the siren acts as a point source of sound.


a. What is the sound intensity at a distance of 6.0m from the siren?


b. What is the power of the sound emitted by the siren?


c. What is the sound intensity level at a distance of 16.0 m from the siren?

1
Expert's answer
2021-12-12T16:40:42-0500

a. For 8 m:


"L_8=90-20\log(8/6)=88\text{ dB}."

b. The power of the sound:


"P=10^{90/10}P_0=0.001\text{ W/m}^2."

c. For 16 m:


"L_{16}=90-20\log(16/6)=81\text{ dB}."